# DESY/PUBDB-2024-01826

### v1

@@ Abstract
Jet flavour tagging is crucial in experimental high-energy physics. A tagging algorithm, DeepJetTransformer, is presented, which exploits a transformer-based neural network that is substantially faster to train.
The DeepJetTransformer network uses information from particle flow-style objects and secondary vertex reconstruction as is standard for $b$- and $c$-jet identification supplemented by additional information, such as reconstructed V0s and $K^{\pm}/\pi^{\pm}$ discrimination, typically not included in tagging algorithms at the LHC. The model is trained as a multiclassifier to identify all quark flavours separately and performs excellently in identifying $b$- and $c$-jets. An $s$-tagging efficiency of $40\%$ can be achieved with a $10\%$ $ud$-jet background efficiency. The impact of including V0s and $K^{\pm}/\pi^{\pm}$ discrimination is presented.
The network is applied on exclusive $Z\to q\bar{q}$ samples to examine the physics potential and is shown to isolate $Z\to s\bar{s}$ events. Assuming all other backgrounds can be efficiently rejected, a $5\sigma$ discovery significance for $Z\to s\bar{s}$ can be achieved with an integrated luminosity of $60~{}\text{nb}^{-1}$, corresponding to less than a second of the FCC-ee run plan at the $Z$ resonance.
@@ 1 Introduction
The Standard Model (SM) of particle physics is one of the most successful scientific theories describing the fundamental particles and their interactions. The last piece of this model, the Higgs boson, was discovered 20121 ; 201230 at the Large Hadron Collider (LHC) Bruning:782076 in $2012$, and the precise study of its properties will remain mostly superficial at the LHC due to high backgrounds and relatively low numbers of Higgs boson events after trigger and selection.
@@
One of the main motivations for proposed future lepton colliders Benedikt:2651299 ; Adolphsen:1601969 ; https://doi.org/10.48550/arxiv.1811.10545 ; Linssen:2012hp is the precise measurement of SM parameters, like precision studies of the hadronic decay of the $Z$ boson and greatly improved sensitivity to the couplings of the Higgs boson to the bottom ($b$) and charm ($c$) quarks and gluons ($g$) European:2720131 ; de_Blas_2020 ; https://doi.org/10.48550/arxiv.1701.02663 . Achieving these objectives requires an efficient reconstruction and identification of the hadronic decays of these particles. The feasibility of studying the decay of the Higgs boson to the strange ($s$) and the light quarks depends on the collider and detector performance and is currently under investigation in the field. It is well established that efficient and accurate jet flavour identification is essential to exploit the maximal physics potential of future collider experiments Azzi2021 ; An_2019 ; https://doi.org/10.48550/arxiv.1310.0763 ; https://doi.org/10.48550/arxiv.2203.07535 ; Abramowicz_2017 .
Jets originating from the $b$ and $c$ quarks contain hadrons with significant lifetimes that travel distances of the order of millimeters from the interaction point before decaying into lighter hadrons. The heavy flavour tagging algorithms used at the Large Electron-Positron collider (LEP) 2004 ; Proriol:1950599 and the Tevatron Abazov_2010 ; Freeman_2013 experiments exploited variables derived from the displaced charged tracks originating from these decayed $B$ (containing $b$ quark) or $D$ (containing $c$ quark) hadrons to distinguish the heavy flavoured jets from the light quark and gluon jets. These charged tracks are commonly clustered to reconstruct the original decay vertices of the $B$ and $D$ hadrons, also called secondary vertices (SVs). The properties of these SVs, like their mass and displacement, can also be used to identify $b$- and $c$-jets.
The understanding and performance of jet flavour tagging at the LHC has steadily been improving and heavily relies on machine learning atlas_run2_ftagging ; CMS:2017wtu , which also inspires flavour tagging algorithms for the FCC-ee Qu:2019gqs ; Bols_2020 . The cleaner environment at lepton colliders and the powerful capabilities of the proposed detectors, such as particle identification (PID), improve the performance of heavy flavour tagging and new analysis techniques, including strange jet tagging, become feasible. Strange jets tend to have a higher kaon multiplicity and a lower number of pions than light jets. Therefore distinguishing $K^{\pm}$ and $\pi^{\pm}$ and reconstructing $K^{0}_{S}$ is crucial for strange jet identification Erdmann_2020 ; Nakai:2020kuu ; Erdmann:2020ovh .
Machine learning (ML) approaches are uniquely suited to classify jet flavours, where training samples are abundant in the form of Monte Carlo (MC) simulation. Still, the underlying dynamics of jet formation and hadronisation are not always well understood. With the advent of Neural Networks (NNs) to jet classification, approaches relying on single physics-motivated variables for jet flavour discrimination were significantly outperformed CMS:2017wtu . Since then, a multitude of architectures and jet representations have found success in discriminating jet flavours, including Dense Neural Networks (DNNs) Luo:2017ncs , Recurrent Neural Networks (RNNs) Guest:2016iqz , Convolutional Neural Networks (CNNs) Komiske:2016rsd ; ATLAS:2017dfg , and Graph Neural Networks (GNNs) Qu:2019gqs ; Mikuni:2020wpr .
Among the most successful of these are Graph-based architectures such as ParticleNet Qu:2019gqs that represent jets as sets of nodes (jet constituents) and edges (some pairwise defined feature, often the difference in a given variable of jet constituents). In particular, networks combining a self-attention mechanism to exploit the relative importance of constructed features, dubbed Transformer Networks, have achieved state-of-the-art performance. Particle Transformer (ParT) Qu:2022mxj combines a graph representation of jets with an attention mechanism. A pure Transformer architecture based on Ref. https://doi.org/10.48550/arxiv.1706.03762 , which is performant and lightweight compared to the state-of-the-art but comparatively computationally expensive ParT, is presented in this work for the task of jet flavour identification at future lepton colliders, using the FCC-ee with the IDEA detector concept as a benchmark Benedikt:2651299 ; Bedeschi:2021bS .
@@
Section 2 summarises the FCC-ee collider, the IDEA detector concept, and the used simulated samples and provides minimal event selection requirements. Section 3 briefly describes the algorithms used to reconstruct displaced decay vertices and their performance. Section 4 introduces the attention mechanism and Transformer models and outlines the description of the input features and the network architecture used for tagging. Finally, the obtained results and the performance of the flavour-tagging algorithm in Z boson signatures are presented in Section 5 and 6, respectively.
@@ 2.1 FCC-ee
The Future Circular Collider (FCC) integrated project FCC:2018byv ; fccfs-midterm aims to build $e^{+}e^{-}$, $pp$, and $ep$ colliders in a $90.7$ km circular tunnel in the Geneva region. FCC-ee Benedikt:2651299 is a proposed lepton collider and the first stage of the FCC integrated project. It is currently planned to run at $4$ different center-of-mass energy modes, starting from around $91.2$ GeV at the Z-pole to $365$ GeV, over the $t\bar{t}$ threshold. The unprecedented luminosities at the FCC-ee uniquely facilitate tests of the SM and, at the same time, present novel challenges in reducing systematic errors. The circular collider design provides the opportunity for multiple interaction points, each of which can host a different detector design. Three such detector concepts Bedeschi:2021bS ; https://doi.org/10.48550/arxiv.1911.12230 ; Francois2022 are currently being studied, of which the IDEA detector concept Bedeschi:2021bS has been used in this study.
@@ 2.3 Event Samples and Jet Reconstruction
The simulated event samples consist of the process $e^{+}e^{-}\rightarrow Z\rightarrow q\bar{q}$, where $q\equiv b,c,(u,d,s)$, at the center-of-mass energy ($\sqrt{s}$) of $91.2$ GeV. Pythia8.303 https://doi.org/10.48550/arxiv.2203.11601 is used for event generation, parton showering, and hadronisation. Delphes de_Favereau_2014 is used for event reconstruction assuming the IDEA detector concept Bedeschi:2021bS ; idea_delphes .
Jet clustering is performed with FastJet-3.3.4 Cacciari_2012 using the exclusive $e^{+}e^{-}~{}k_{\text{T}}$ algorithm Catani:1991hj . Other jet clustering algorithms like the anti-$k_{\text{T}}$ algorithm Cacciari_2008 and the generalised $e^{+}e^{-}~{}k_{\text{T}}$, also referred to as the inclusive $e^{+}e^{-}~{}k_{\text{T}}$, algorithm Cacciari_2012 were also considered. The exclusive $e^{+}e^{-}~{}k_{\text{T}}$ algorithm clustered jets were observed to satisfy the requirements of this study and signature the best.
@@ 3 Vertex Reconstruction
Vertex reconstruction is essential to find the primary interaction vertex and the secondary decay vertices of the long-lived $B$, $D$, and $S$ (containing $s$ quark) hadrons. It helps improve the $b$- and $c$-tagging performance and aids in $s$-tagging. Charged tracks can be fitted to reconstruct the primary and the displaced secondary vertices. These displaced vertices can either be the decay vertices of $B$ and $D$ hadrons (SVs) or those of the long-lived $S$ hadrons, like $K^{0}_{S}$ or $\Lambda^{0}$, also known as V0s, which are particles that decay into a pair of oppositely charged tracks. The properties of these SVs, such as their masses, displacements, and charged track multiplicities, can be used to identify the decaying hadrons and, in effect, the jet flavour. The SVs can even be used to reconstruct the entire hadronic decay chain. Similarly, reconstructing and identifying the V0 vertices can be used to identify $s$-jets, as $K^{0}_{S}$ and $\Lambda^{0}$ are the leading particles in some $s$-jets. Distinguishing V0s from SVs also helps to reduce the misidentification of some $b$- and $c$-jets as $s$-jets. The vertex reconstruction in this study has been performed using an implementation of the vertexing module of the LCFIPlus framework Suehara_2016 ; Gautam:2839048 as implemented in FCCAnalyses fccana , the FCC software framework, using a $\chi^{2}$-based vertex fitter vertex_fitter .
The properties of these SVs and V0s, along with more variables, are used as input to train the neural network tagger described in Section 4.
(a) $K_{S}^{0}$
(b) $\Lambda^{0}$
Figure 1: Invariant mass distribution of reconstructed V0 vertices in jets from $e^{+}e^{-}\rightarrow Z\rightarrow q^{\prime}\bar{q^{\prime}}$ events at $\sqrt{s}=91.2$ GeV, where $q^{\prime}\equiv u,d,s,c,b$.

### v4

@@ Abstract
Jet flavour tagging is crucial in experimental high-energy physics. A tagging algorithm, DeepJetTransformer, is presented, which exploits a transformer-based neural network that is substantially faster to train than state-of-the-art graph neural networks.
The DeepJetTransformer algorithm uses information from particle flow-style objects and secondary vertex reconstruction for $b$- and $c$-jet identification, supplemented by additional information that is not always included in tagging algorithms at the LHC, such as reconstructed $K_{S}^{0}$ and $\Lambda^{0}$ and $K^{\pm}/\pi^{\pm}$ discrimination. The model is trained as a multiclassifier to identify all quark flavours separately and performs excellently in identifying $b$- and $c$-jets. An $s$-tagging efficiency of $40\%$ can be achieved with a $10\%$ $ud$-jet background efficiency. The performance improvement achieved by including $K_{S}^{0}$ and $\Lambda^{0}$ reconstruction and $K^{\pm}/\pi^{\pm}$ discrimination is presented.
The algorithm is applied on exclusive $Z\to q\bar{q}$ samples to examine the physics potential and is shown to isolate $Z\to s\bar{s}$ events. Assuming all non-$Z\to q\bar{q}$ backgrounds can be efficiently rejected, a $5\sigma$ discovery significance for $Z\to s\bar{s}$ can be achieved with an integrated luminosity of $60~{}\text{nb}^{-1}$ of $e^{+}e^{-}$ collisions at $\sqrt{s}=91.2~{}\mathrm{GeV}$, corresponding to less than a second of the FCC-ee run plan at the $Z$ boson resonance.
@@ 1 Introduction
The Standard Model (SM) of particle physics Glashow:1959wxa ; Salam:1959zz ; Weinberg1967 ; tHooft1972 is one of the most successful scientific theories describing the fundamental particles and their interactions. The last piece of this model, the Higgs boson, was discovered 20121 ; 201230 at the Large Hadron Collider (LHC) Bruning:782076 in $2012$, and the precise study of its properties will remain mostly superficial at the LHC due to high irreducible backgrounds from other SM processes while isolating Higgs boson events.
@@
One of the main motivations for proposed future lepton colliders Benedikt:2651299 ; Adolphsen:1601969 ; CEPCStudyGroup:2018ghi ; Linssen:2012hp is the precise measurement of SM parameters, like precision studies of the hadronic decay of the $Z$ boson and greatly improved sensitivity to the couplings of the Higgs boson to the bottom ($b$) and charm ($c$) quarks and gluons ($g$) European:2720131 ; de_Blas_2020 ; dEnterria:2017dac . Achieving these objectives requires an efficient reconstruction and identification of the hadronic decays of these particles. The feasibility of studying the decay of the Higgs boson to the strange ($s$), up ($u$), and down ($d$) quarks depends on the collider and detector performance and is currently under investigation in the field. It is well established that efficient and accurate jet flavour identification is essential to exploit the maximal physics potential of future collider experiments Azzi2021 ; An_2019 ; Asner:2013psa ; Albert:2022mpk ; Abramowicz_2017 .
@@
The state-of-the art is shortly reviewed in the rest of Section 1. Section 2 summarises the FCC-ee collider, the IDEA detector concept, and the used simulated samples and provides minimal event selection requirements. Section 3 briefly describes the algorithms used to reconstruct displaced decay vertices and their performance. Section 4 introduces the attention mechanism and Transformer models and outlines the description of the input features and the network architecture used for tagging. Finally, the obtained results and the performance of the flavour-tagging algorithm in $Z$ boson signatures are presented in Section 5 and 6, respectively.
1.1 Review of Jet Flavour Tagging
Jets originating from the $b$ and $c$ quarks contain hadrons with significant lifetimes that travel distances of the order of millimeters from the interaction point before decaying into lighter hadrons. The heavy flavour tagging algorithms used at the Large Electron-Positron collider (LEP) 2004 ; Proriol:1950599 and the Tevatron Abazov_2010 ; Freeman_2013 experiments exploited variables derived from the displaced charged tracks originating from these decayed $b$ or $c$ hadrons to distinguish the heavy flavoured jets from $s$, $u$, $d$ quark and gluon jets. These charged tracks are commonly clustered to reconstruct the original decay vertices of the $b$ and $c$ hadrons, also called secondary vertices (SVs). The properties of these SVs, like their mass and displacement, are some of the most important inputs used to identify $b$- and $c$-jets.
The understanding and performance of jet flavour tagging at the LHC has steadily been improving and heavily relies on machine learning (ML) atlas_run2_ftagging ; CMS:2017wtu , which also inspires flavour tagging algorithms for the FCC-ee Qu:2019gqs ; Bols_2020 .
ML approaches are uniquely suited to classify jet flavours, where training samples are abundant in the form of Monte Carlo (MC) simulation. Still, the underlying dynamics of jet formation and hadronisation are not always well understood. With the advent of ML techniques, including Neural Networks (NNs) and Boosted Decision Trees (BDTs), approaches relying on single physics-motivated variables for jet flavour discrimination were significantly outperformed CMS:2017wtu ; ATLAS:2015thz ; ATLAS:2019bwq ; Mondal_2024 . Since then, a multitude of architectures and jet representations have found success in discriminating jet flavours, including Dense Neural Networks (DNNs) Luo:2017ncs , Recurrent Neural Networks (RNNs) Guest:2016iqz , Convolutional Neural Networks (CNNs) Komiske:2016rsd ; ATLAS:2017dfg , and Graph Neural Networks (GNNs) Qu:2019gqs ; Mikuni:2020wpr ; ATLAS:2022rkn .
Among the most successful of these are Graph-based architectures such as ParticleNet Qu:2019gqs that represent jets as sets of nodes (jet constituents) and edges (some pairwise defined feature, often the difference in a given variable of jet constituents). In particular, networks combining a self-attention mechanism vaswani2023attentionneed to exploit the relative importance of constructed features, dubbed Transformer Networks, have achieved state-of-the-art performance in the task of jet flavour tagging Mondal_2024 ; duperrin2023flavourtagginggraphneural ; ATLAS:2023ber ; Qu:2022mxj . Particle Transformer (ParT) Qu:2022mxj combines a graph representation of jets with an attention mechanism. In this work, a pure Transformer architecture, DeepJetTransformer, similar to the ParT (plain) variant introduced in Ref. Qu:2022mxj , is presented for the task of jet flavour identification at future lepton colliders, using the FCC-ee with the IDEA detector concept as a benchmark Benedikt:2651299 ; Bedeschi:2021bS . DeepJetTransformer is relatively lightweight and requires much less computational time compared to graph-based architectures Bedeschi_2022 ; Qu:2022mxj ; 10363595 , yet achieves comparable tagging performance.
1.2 The Z boson at the FCC-ee
Future colliders with a dedicated $Z$ boson run, like FCC-ee, will improve the precision of all these measurements and make the $s$ quark, and potentially the $u$, and $d$ quarks, accessible. Individual measurements of the quark vector and axial couplings should be possible via their forward-backward asymmetries, corresponding partial decay widths of the $Z$ boson, and the precise knowledge of $A_{e}$, the asymmetry parameter of the $e^{-}e^{+}$ pair. The experimental systematic uncertainties corresponding to these measurements are also expected to drastically improve due to better detector designs Benedikt:2651299 .
1.3 Strange Jet Tagging
The discrimination of $s$-jets is widely regarded as one of the most challenging types of jet discrimination. Thus, it has received considerably less attention than its heavy-flavour counterparts, or indeed gluon discrimination. At the core of the problem is the fact that unlike in the discrimination of quarks vs gluons, which relies heavily on properties following from their differing colour factors $C_{F}=4/3$ vs $C_{A}=3$, or heavy flavour tagging, which relies on displaced vertices of $b/c$ hadrons, strange quarks are treated identically to down quarks by QCD and Electroweak theory in the massless limit prior to their decay. Discriminating strange and down jets is particularly challenging due to the same fractional charge of the initiating quarks. In practice, however, strange hadrons carry a larger fraction of the total scalar momentum of strange jets, compared to hadrons consisting of up and down ($ud$) quarks. The total scalar momentum is obtained by summing over the scalar momentum of all jet constituents. This idea was also explored in the context of hadron colliders Nakai:2020kuu . Strange jets tend to have a higher kaon multiplicity and a lower number of pions than $u$- and $d$-jets. Therefore distinguishing $K^{\pm}$ and $\pi^{\pm}$ and reconstructing $K^{0}_{S}$ is crucial for strange jet identification Erdmann_2020 ; Nakai:2020kuu ; Erdmann:2020ovh .
SLD SLD:1984aa tagged $Z\to s\bar{s}$ events by looking for the absence of reconstructed $b$ and $c$ hadrons and the presence of $K^{\pm}$ or $K^{0}_{S}$ PhysRevLett.85.5059 . Particle identification (PID) was performed at SLD, as at DELPHI delphi_loi , with a RICH detector. At most other detectors, energy loss ($dE/dx$) was used for PID opal_loi ; l3_loi , with the addition of timing at ALEPH aleph_loi . The detector concepts at the FCC-ee foresee the use of techniques like energy loss ($dE/dx$) Lippmann2012 , ionisation cluster counting ($dN/dx$) Cataldi:1996mz , time-of-flight Klempt:1999tx , and compact-Ring Imaging CHerenkov (RICH) detectors.
Tagging strange jets at future colliders has been explored as a probe to perform precision measurements in the Higgs sector DuarteCampderros2020 ; Albert:2022mpk , and the impact of using $dN/dx$ and time-of-flight on strange tagging performance for jets originating from Higgs boson decay was studied using a graph neural network Bedeschi_2022 . In this work, DeepJetTransformer is used to isolate $Z\to s\bar{s}$ events from the exclusive hadronic decays of the $Z$ boson in the FCC-ee environment. The excess momentum carried by strange hadrons is exploited, firstly by including V0 variables and secondly through $K^{\pm}/\pi^{\pm}$ discrimination. The cleaner environment at lepton colliders and the powerful PID capabilities of the proposed detectors facilitate making strange jet tagging feasible.
@@ 2.1 FCC-ee
The Future Circular Collider (FCC) integrated project FCC:2018byv ; fccfs-midterm aims to build $e^{+}e^{-}$, $pp$, and $ep$ colliders in a $90.7$ km circular tunnel in the Geneva region. FCC-ee Benedikt:2651299 is a proposed $e^{+}e^{-}$ collider and the first stage of the FCC integrated project. It is currently planned to run at four different center-of-mass energy modes, starting from around $91.2$ GeV at the Z-pole to $365$ GeV, over the $t\bar{t}$ threshold. The unprecedented luminosities at the FCC-ee uniquely facilitate tests of the SM and, at the same time, present novel challenges in reducing systematic errors. The circular collider design provides the opportunity for four interaction points, each of which can host a different detector design. Such detector concepts Bedeschi:2021bS ; https://doi.org/10.48550/arxiv.1911.12230 ; Francois2022 are currently being studied, of which the IDEA detector concept Bedeschi:2021bS has been used in this study.
@@ 2.2 IDEA Detector Concept
A fast simulation of the IDEA detector concept idea_delphes has been implemented in Delphes de_Favereau_2014 and used for the simulation of the samples used in this work. A spherical coordinate system is used with its origin at the center of the detector system and the positive $z$ axis in the direction of travel of the incoming electron. The polar angle, $\theta$, is defined as the angle between the radial line and the positive $z$ axis and the azimuthal angle, $\phi$, is defined as the angle of rotation of the radial line around the positive $z$ axis.
@@ 2.3 Event Samples and Jet Reconstruction
The simulated event samples used for training and evaluation consist of the process $e^{+}e^{-}\rightarrow Z\rightarrow q\bar{q}$, where $q\equiv b,c,(u,d,s)$, at the center-of-mass energy ($\sqrt{s}$) of $91.2$ GeV. Pythia8.303 https://doi.org/10.48550/arxiv.2203.11601 is used for event generation, parton showering, and hadronisation. Delphes de_Favereau_2014 is used for event reconstruction assuming the IDEA detector concept Bedeschi:2021bS ; idea_delphes . A tracking efficiency of $99.7\%$ is assumed for electrons, muons and charged hadrons with 3-momentum magnitude $|p|>0.5$ GeV that lie within acceptance. This efficiency is reduced to $65\%$ ($4\%$) for $0.30.5>|p|>$ 0.3 GeV ($|p|<0.3$ GeV). Fake tracks are not considered.
Jet clustering is performed on the particle flow-style objects reconstructed by Delphes with FastJet-3.3.4 Cacciari_2012 using the exclusive $e^{+}e^{-}~{}k_{\text{T}}$ algorithm Catani:1991hj . Other jet clustering algorithms like the anti-$k_{\text{T}}$ algorithm Cacciari_2008 and the generalised $e^{+}e^{-}~{}k_{\text{T}}$, also referred to as the inclusive $e^{+}e^{-}~{}k_{\text{T}}$, algorithm Cacciari_2012 were also considered. The exclusive $e^{+}e^{-}~{}k_{\text{T}}$ algorithm, which creates irregularly shaped jets and, in this study, requires exactly two jets, is very robust against gluon emissions and gluon splitting. Since the exclusive $e^{+}e^{-}~{}k_{\text{T}}$ algorithm clustered jets include all reconstructed final particles, they were observed to satisfy the requirements of this study by most accurately reproducing the $Z$ boson reconstructed invariant mass signature. No additional selections were applied to the samples for training and evaluation of the jet flavour tagger.
In this study, the jets are assigned an MC flavour as the flavour of the quarks to which the $Z$ boson decays. Besides simplicity, this has the added benefit that other studies for future facilities use the same definition.
A separate set of event samples was generated with the process $e^{+}e^{-}\rightarrow Z(\rightarrow\nu\nu)H\rightarrow q\bar{q}$, where $q\equiv b,c,(u,d,s)$, at $\sqrt{s}$ of $240$ GeV. The same reconstruction and jet clustering were applied as for the samples at the $Z$ resonance. Training and evaluation of DeepJetTransformer were performed with these samples for comparison with other taggers.
@@ 3 Vertex Reconstruction
Vertex reconstruction is essential to find the primary interaction vertex and the secondary decay vertices of the long-lived $b$, $c$, and $s$ hadrons. It helps improve the $b$- and $c$-tagging performance and aids in $s$-tagging. Charged tracks can be fitted to reconstruct the primary and the displaced secondary vertices. These displaced vertices can either be the decay vertices of $b$ and $c$ hadrons (SVs) or those of the long-lived hadrons containing $s$ quarks, like $K^{0}_{S}$ or $\Lambda^{0}$, commonly referred to as V0s, which are particles that decay into a pair of oppositely charged tracks. All displaced vertices except V0s are referred to as SVs. The properties of SVs and V0s, such as their masses, displacements, and charged track multiplicities, can be used to identify the decaying hadrons and, in effect, the jet flavour. The SVs can even be used to reconstruct the entire hadronic decay chain. Similarly, reconstructing and identifying the V0 vertices can be used to identify $s$-jets, as $K^{0}_{S}$ and $\Lambda^{0}$ are the particles carrying most of the momentum of some $s$-jets Albert:2022mpk . Distinguishing V0s from SVs also helps to reduce the misidentification of some $b$- and $c$-jets as $s$-jets.
The vertex reconstruction in this study has been performed using an implementation of the vertexing module of the LCFIPlus framework Suehara_2016 ; Gautam:2839048 . It has been implemented in FCCAnalyses fccana , the FCC software framework, using a $\chi^{2}$-based vertex fitter vertex_fitter . The constraints and parameters have been kept the same as in Ref. Suehara_2016 . The algorithm first identifies the tracks forming V0s. Unlike in standard vertex reconstruction algorithms, the V0s are not discarded but stored and assigned a particle ID based on the set of constraints that they pass, summarised in Table 1. The tracks originating from the primary vertex or V0 candidates are not considered while reconstructing SVs.
The properties of the SVs and V0s, along with more variables, are used as input to train the neural network tagger described in Section 4.
(a) Reconstructed $K_{S}^{0}$
(b) V0 multiplicities
Figure 1: Performance of V0 reconstruction. (a) Invariant mass distribution of reconstructed $K_{S}^{0}$ vertices. The quoted mass is the mean and the error on the mean of the distribution. (b) The reconstructed V0 multiplicity in jets from $e^{+}e^{-}\rightarrow Z\rightarrow q\bar{q}$ events at $\sqrt{s}=91.2$ GeV, where $q\equiv u,d,s,c,b$. The distributions for $b$- and $s$-jets overlap almost perfectly.